# The Second Workshop on GenAI for Health: Potential, Trust, and Policy Compliance

### v2

@@ Abstract
Medical question answering requires advanced reasoning that integrates domain knowledge with logical inference. However, existing large language models (LLMs) often generate reasoning chains that lack factual accuracy and clinical reliability. We propose Ranked Preference Reinforcement Optimization (RPRO), a novel framework that uniquely combines reinforcement learning with preference-driven reasoning refinement to enhance clinical chain-of-thought (CoT) performance. RPRO differentiates itself from prior approaches by employing task-adaptive reasoning templates and a probabilistic evaluation mechanism that aligns outputs with established clinical workflows, while automatically identifying and correcting low-quality reasoning chains. Unlike traditional pairwise preference methods, RPRO introduces a groupwise ranking optimization based on the Bradley–Terry model and incorporates KL-divergence regularization for stable training. Experiments on PubMedQA and MedQA-USMLE show consistent improvements over strong baselines. Remarkably, our 1.1B parameter model outperforms much larger 7B–13B models, including medical-specialized variants. These findings demonstrate that combining preference optimization with quality-driven refinement offers a scalable and effective approach to building more reliable, clinically grounded medical LLMs.
@@ 1 Introduction
Recent advances in large language models (LLMs) signal significant promise for medical applications, especially in clinical reasoning and diagnostic support systems singhal2023large; thirunavukarasu2023large. However, deploying AI in medicine requires exceptional reliability, making factual accuracy and thorough coverage of the data paramount. Medical reasoning introduces challenges distinct from those in general domains: clinical decisions require a comprehensive consideration of relevant factors, strict factual accuracy, and the avoidance of redundant information that could obscure critical insights.
Current approaches to medical reasoning primarily rely on chain-of-thought (CoT) prompting techniques wei2022chain; kojima2022large, which have shown effectiveness in improving reasoning capabilities across various domains. However, these methods typically generate single reasoning chains without systematic quality assessment or iterative refinement mechanisms. Moreover, existing evaluation frameworks often rely on overly simplistic binary metrics that fail to capture the nuanced quality required for medical reasoning or require extensive human annotation, limiting their suitability for automated quality control where domain expertise is essential.
Recent work in preference learning and reinforcement learning from human feedback (RLHF) ouyang2022training; christiano2017deep shows promise in aligning AI systems with human preferences. Direct Preference Optimization (DPO) rafailov2023direct has shown promising results by learning from preference data. However, these methods typically rely on pairwise comparisons, which fail to capture all the ranking information in human preference datasets. Many real-world scenarios involve humans ranking multiple candidates, not just providing binary choices. For example, when they evaluate reasoning chains, mathematical solutions, or creative content, evaluators assign rankings that reflect nuanced quality differences. Reducing this rich data to pairwise comparisons can lose valuable training signals and lead to suboptimal solutions.
Specifically, these approaches face significant challenges in medical domains. First, the requirement for extensive expert annotation is prohibitively expensive. Additionally, preference judgments often lack the granularity needed for medical reasoning assessment. Furthermore, generic preference models fail to capture the domain-specific quality dimensions that are critical in healthcare applications.
@@ 2.1 Medical LLMs: QA, Diagnosis, and EHR
Large language models (LLMs) have been used for medical tasks like knowledge-based question answering, diagnostic reasoning on clinical vignettes, and patient-specific reasoning over electronic health records (EHR) (jin2019pubmedqa; jin2021medqa; shi2024ehragent; zakka2024almanac). Domain-structured prompts and chain-of-thought (CoT) approaches consistently improve auditability and reduce unsupported claims by clearly separating decomposition and background from justification or diagnosis (wei2022cot; singhal2023medpalm; singhal2023medpalm2). In patient-specific contexts, retrieval-augmented generation (RAG) and citation-style evidence tracing help ensure faithfulness to chart data and guidelines lewis2020rag; shi2024ehragent; zakka2024almanac. Recent work extends beyond answer accuracy, emphasizing multidimensional evaluation such as coverage of relevant findings and alignment with context, thereby better matching clinical expectations for completeness and correctness. In this work, we use domain-specific CoT templates (QA and diagnosis), automatically score rationales on coverage, factuality, and redundancy with an LLM judge, and make light score-guided revisions.
@@ 2.2 Reinforcement-Learned Medical LLMs
Reinforcement learning from human feedback (RLHF) and related preference-based methods have become central to aligning LLM behavior beyond supervised fine-tuning christiano2017deep; ziegler2019fine; ouyang2022training. In reasoning-heavy domains, variants that regularize to a reference policy while optimizing rewards derived from preference ordering or task-specific outcomes have shown improved stability and faithfulness.
In the medical domain, reinforcement-learned LLMs have been applied to tasks such as clinical question answering and diagnostic reasoning singhal2023towards; nori2023capabilities. These methods typically employ reinforcement learning techniques such as policy optimization with preference feedback christiano2017deep; ouyang2022training to guide models toward multi-step clinical reasoning, integration of biomedical knowledge, and reliable elimination of alternative diagnoses. Nevertheless, prior studies also emphasize persistent challenges in maintaining stability and factual grounding in high-stakes clinical scenarios, indicating that preference-optimized training plays an important role in reducing hallucination risks singhal2023towards.
@@ 2.3 Ranking-based Optimization in Clinical Reasoning
Recent studies have expanded preference alignment for LLMs from pairwise to listwise settings. Most methods focus on pairwise comparisons. Yet, listwise approaches have shown effectiveness in other domains liu2024lipo; pesaran2025lpoi; cai2025kpo. For example, ALMupQA by yang2024aligning integrates multi-perspective ranking alignment for code QA. IRPO, introduced by wu2025context, is a framework that incorporates graded relevance and positional importance. These works highlight the impact of ranking-based feedback on aligning LLM outputs with diverse user expectations. However, medical multiple-choice QA often needs more sophisticated explanatory discourse and clinical reasoning. In this work, we evaluate our framework on multiple medical reasoning benchmarks. We use a probabilistic refinement mechanism, generate preference data with RPRO, and combine linear rewards for ranked candidate optimization.
@@ 3.3 Task-Adaptive Chain-of-Thought Generation
For $\tau=\tau_{QA}$, the components include four clinical factor $s^{(1)}:\ \text{Question Decomposition}$, $s^{(2)}:\ \text{Background Knowledge}$, $s^{(3)}:\ \text{Logical Connection}$ and $s^{(4)}:\ \text{Final Justification}$ to determine generated sequence medical QA chain is correct. Similarly, the $\tau=\tau_{Diag}$ focus on diagnosis task include $t^{(1)}:\ \text{Case Summary}$, $t^{(2)}:\ \text{Clinical Significance}$, $t^{(3)}:\ \text{Differential Diagnosis}$ and $t^{(4)}:\ \text{Most Likely Diagnosis}$. Those criteria can better align the reasoning process with standard diagnostic procedures, thereby tailoring the CoT representation to the requirements of clinical decision-making. Thus, the unified formulation $c_{k}=(u^{(1)}_{k},u^{(2)}_{k},u^{(3)}_{k},u^{(4)}_{k})$ provides a task-agnostic representation of CoT candidates, while retaining the ability to specialize the semantic meaning of each reasoning step according to the task label $\tau$. After generation, the set of five candidates is reduced to $M=4$ by retaining the top four according to subsequent evaluation and discarding the lowest-ranked one.
@@ 3.4 Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning zhu2024llms; wei2022chain; kojima2022large. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table 3 presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section 3.3. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ 4.1 Ranking Preference Optimization Algorithm
where $|c_{j}|$ denotes the sequence length of the CoT, and $c_{j,t}$ represents the token at position $t$ in the reasoning chain. Following the Bradley-Terry model bradley1952rank, the probability that CoT sequence $c_{i}$ is preferred over sequence $c_{j}$ is:
@@ 4.2 Linear Rewards From Rank
GRPO is one of the widely used algorithms that employs entropy regularization to stabilize policy gradient–based learning processes. However, it still exhibits sensitivity to parameter tuning, which can lead to unstable convergence performance or even divergence during training leahy2022convergence; schulman2017proximal; mnih2016asynchronous. To translate the complete ranking of candidate reasoning chains into effective training signals, we adopt a linear reward shaping scheme. We then define the advantage value $a_{j}$ of candidate $c_{j}$ as:
@@ 5.1 Experimental Setting
Datasets. We evaluate our approach on two benchmark medical QA datasets. PubMedQA jin2019pubmedqa contains biomedical research questions derived from PubMed abstracts, requiring yes/no/maybe answers based on scientific evidence. MedQA-USMLE jin2021disease consists of multiple-choice diagnostic questions from the United States Medical Licensing Examination, testing clinical reasoning and diagnostic capabilities. These datasets represent complementary aspects of medical reasoning: research-oriented knowledge synthesis and clinical diagnostic decision-making.
Implementation Details. Our base model is LLaMA 1.1B, chosen for computational efficiency while maintaining competitive performance. We generate five CoT reasoning variants per question and select the four highest-ranking candidates based on their acceptance probabilities for RPRO training. The acceptance threshold $\tau$ is set to 0.6 based on preliminary experiments. All models are trained with AdamW optimizer using a learning rate of 5e-5, batch size of 16, and 3 training epochs. Training is performed on NVIDIA A100 GPUs with mixed precision.
@@ 5.2 Baselines
We compare against several established language models across different scales and medical specializations. General-purpose models include Gemma 2B and 7B mesnard2024gemma, and Mistral 7B jiang2023mistral. Medical-specialized models include MedAlpaca 7B and 13B han2023medalpaca, fine-tuned specifically for medical tasks, and BioMistral 7B labrak2024biomistral, trained on biomedical corpora. These baselines represent the current state-of-the-art in both general and domain-specific language modeling for medical applications.
@@ 5.4 Evaluation Metrics
5.5 Quantitative Results
Table 1: Comparison of different models on PubMedQA and MedQA-USMLE. Best results are in bold.
Models Method PubMedQA MedQA-USMLE
ACC (%) Macro F1 (%) ACC (%) Macro F1 (%)
Gemma 2B - 51.03 52.61 17.78 9.21
Gemma 7B - 58.08 48.02 20.37 10.18
Mistral 7B - 51.43 47.11 22.66 11.78
MedAlpaca 7B - 44.83 42.41 24.01 16.37
BioMistral 7B - 38.89 37.98 19.89 7.47
MedAlpaca 13B - 52.61 46.16 23.15 22.42
Our Method (LLaMA 1.1B, RPRO) RPRO 61.95 50.88 27.92 20.41
Table 2: Ablation study comparing different training methods and refinement strategies on PubMedQA and MedQA-USMLE. Best results are in bold.
@@ 5.5 Quantitative Results
SFT (with Refinement) 57.46 41.37 24.81 13.52
DPO (with Refinement) 60.90 41.92 25.83 15.65
RPRO (No Refinement) 61.43 38.11 23.33 13.27
Our Method: RPRO (with Refinement) 61.95 50.88 27.92 20.41

### v3

# RPRO: Ranked Preference Reinforcement Optimization for Enhancing Medical QA and Diagnostic Reasoning

@@ Abstract
Medical question answering requires advanced reasoning that integrates domain knowledge with logical inference. However, existing large language models (LLMs) often generate reasoning chains that lack factual accuracy and clinical reliability. We propose Ranked Preference Reinforcement Optimization (RPRO), a novel framework that combines reinforcement learning with preference-driven reasoning refinement to enhance clinical chain-of-thought (CoT) performance. RPRO distinguishes itself from prior approaches by employing task-adaptive reasoning templates and a probabilistic evaluation mechanism that aligns model outputs with established clinical workflows, while automatically identifying and correcting low-quality reasoning chains. Unlike traditional pairwise preference methods, RPRO introduces a groupwise ranking optimization based on the Bradley–Terry model and incorporates KL-divergence regularization for stable training. Experiments on PubMedQA, MedQA-USMLE, and a real-world clinical dataset from Far Eastern Memorial Hospital (FEMH) demonstrate consistent improvements over strong baselines. Remarkably, our 2B-parameter model outperforms much larger 7B–20B models, including medical-specialized variants. These findings demonstrate that combining preference optimization with quality-driven refinement provides a scalable and clinically grounded approach to building more reliable medical LLMs.
keywords:
Electronic Health Records, Medical QA, Diagnostic Reasoning; Preference Optimization, Reinforcement Learning, Chain-of-Thought
[…]
@@ 1 Introduction
Recent advances in large language models (LLMs) signal significant promise for medical applications, especially in clinical reasoning and diagnostic support systems [singhal2023large, thirunavukarasu2023large, si2023evaluating, gilson2023does]. However, deploying AI in medicine requires exceptional reliability, making factual accuracy and thorough coverage of the data paramount. Medical reasoning introduces challenges distinct from those in general domains: clinical decisions require a comprehensive consideration of relevant factors, strict factual accuracy, and the avoidance of redundant information that could obscure critical insights.
Current approaches to medical reasoning primarily rely on chain-of-thought (CoT) prompting techniques [wei2022chain, kojima2022large, wang2023selfconsistency], which have shown effectiveness in improving reasoning capabilities across various domains. However, these methods typically generate single reasoning chains without systematic quality assessment or iterative refinement mechanisms. Moreover, existing evaluation frameworks often rely on overly simplistic binary metrics that fail to capture the nuanced quality required for medical reasoning or require extensive human annotation, limiting their suitability for automated quality control where domain expertise is essential.
Recent work in preference learning and reinforcement learning from human feedback (RLHF) [ouyang2022training, christiano2017deep, bai2022training] shows promise in aligning AI systems with human preferences. Direct Preference Optimization (DPO) [rafailov2023direct] has shown promising results by learning from preference data. However, these methods typically rely on pairwise comparisons, which fail to capture all the ranking information in human preference datasets. Many real-world scenarios involve humans ranking multiple candidates, not just providing binary choices. For example, when they evaluate reasoning chains, mathematical solutions, or creative content, evaluators assign rankings that reflect nuanced quality differences. Reducing this rich data to pairwise comparisons can lose valuable training signals and lead to suboptimal solutions.
Specifically, these approaches face significant challenges in medical domains. First, the requirement for extensive expert annotation is prohibitively expensive [talmor2020olmpics, joshi2023doctorsai]. Additionally, preference judgments often lack the granularity needed for medical reasoning assessment [lin2022human]. Furthermore, generic preference models fail to capture the domain-specific quality dimensions that are critical in healthcare applications [nori2023capabilities].
@@ 2.1 Medical LLMs: QA, Diagnosis, and EHR
Large language models (LLMs) have been used for medical tasks like knowledge-based question answering, diagnostic reasoning on clinical vignettes, and patient-specific reasoning over electronic health records (EHR) [jin2019pubmedqa, jin2021medqa, shi2024ehragent, zakka2024almanac]. Domain-structured prompts and chain-of-thought (CoT) approaches consistently improve auditability and reduce unsupported claims by clearly separating decomposition and background from justification or diagnosis [wei2022cot, singhal2023medpalm, singhal2023medpalm2]. In patient-specific contexts, retrieval-augmented generation (RAG) and citation-style evidence tracing help ensure faithfulness to chart data and guidelines [lewis2020rag, shi2024ehragent, zakka2024almanac]. Recent work extends beyond answer accuracy, emphasizing multidimensional evaluation such as coverage of relevant findings and alignment with context, thereby better matching clinical expectations for completeness and correctness. In this work, we use domain-specific CoT templates (QA and diagnosis), automatically score rationales on coverage, factuality, and redundancy with an LLM judge, and make light score-guided revisions.
@@ 2.2 Reinforcement-Learned Medical LLMs
Reinforcement learning from human feedback (RLHF) and related preference-based methods have become central to aligning LLM behavior beyond supervised fine-tuning [christiano2017deep, ziegler2019fine, ouyang2022training]. In reasoning-heavy domains, variants that regularize to a reference policy while optimizing rewards derived from preference ordering or task-specific outcomes have shown improved stability and faithfulness.
In the medical domain, reinforcement-learned LLMs have been applied to tasks such as clinical question answering and diagnostic reasoning [singhal2023towards, nori2023capabilities]. These methods typically employ reinforcement learning techniques such as policy optimization with preference feedback [christiano2017deep, ouyang2022training] to guide models toward multi-step clinical reasoning, integration of biomedical knowledge, and reliable elimination of alternative diagnoses. Nevertheless, prior studies also emphasize persistent challenges in maintaining stability and factual grounding in high-stakes clinical scenarios, indicating that preference-optimized training plays an important role in reducing hallucination risks [singhal2023towards].
@@ 2.3 Ranking-based Optimization in Clinical Reasoning
Recent studies have expanded preference alignment for LLMs from pairwise to listwise settings. Most methods focus on pairwise comparisons. Yet, listwise approaches have shown effectiveness in other domains [liu2024lipo, pesaran2025lpoi, cai2025kpo]. For example, ALMupQA by [yang2024aligning] integrates multi-perspective ranking alignment for code QA. IRPO, introduced by [wu2025context], is a framework that incorporates graded relevance and positional importance. These works highlight the impact of ranking-based feedback on aligning LLM outputs with diverse user expectations. However, medical multiple-choice QA often needs more sophisticated explanatory discourse and clinical reasoning. In this work, we evaluate our framework on multiple medical reasoning benchmarks. We use a probabilistic refinement mechanism, generate preference data with RPRO, and combine linear rewards for ranked candidate optimization.
@@ 3.3 Task-Adaptive Chain-of-Thought Generation
For $\tau=\tau_{QA}$, the components include four clinical factors:
$\begin{array}[]{ll}s^{(1)}:&\text{Question Decomposition},\\ s^{(2)}:&\text{Background Knowledge},\\ s^{(3)}:&\text{Logical Connection},\\ s^{(4)}:&\text{Final Justification.}\end{array}$
to determine generated sequence medical QA chain is correct. Similarly, for $\tau=\tau_{Diag}$, which focuses on the diagnosis task, the components are defined as:
$\begin{array}[]{ll}t^{(1)}:&\text{Case Summary},\\ t^{(2)}:&\text{Clinical Significance},\\ t^{(3)}:&\text{Differential Diagnosis},\\ t^{(4)}:&\text{Most Likely Diagnosis.}\end{array}$
. Those criteria can better align the reasoning process with standard diagnostic procedures, thereby tailoring the CoT representation to the requirements of clinical decision-making. Thus, the unified formulation $c_{k}=(u^{(1)}_{k},u^{(2)}_{k},u^{(3)}_{k},u^{(4)}_{k})$ provides a task-agnostic representation of CoT candidates, while retaining the ability to specialize the semantic meaning of each reasoning step according to the task label $\tau$. After generation, the set of five candidates is reduced to $M=4$ by retaining the top four according to subsequent evaluation and discarding the lowest-ranked one.
@@ 3.4 Medical Prompt Instruction and Reasoning
Previous research shows that effective prompt template instructions can improve model performance on CoT reasoning [zhu2024llms, wei2022chain, kojima2022large]. We use structured reasoning templates based on established clinical and biomedical reasoning patterns. Table 5 presents a general medical QA template that applies a four-step reasoning process, using binary classification at each step according to the validation criteria from Section 3.3. Specifically, background knowledge retrieval gathers relevant biomedical facts and principles from medical literature. Logical connection links the retrieved knowledge to the question context. The justification step synthesizes the reasoning to support the conclusion.
@@ 4.1 Ranking Preference Optimization Algorithm
where $|c_{j}|$ denotes the sequence length of the CoT, and $c_{j,t}$ represents the token at position $t$ in the reasoning chain. Following the Bradley-Terry model [bradley1952rank], the probability that CoT sequence $c_{i}$ is preferred over sequence $c_{j}$ is:
@@ 4.2 Linear Rewards From Rank
GRPO is one of the widely used algorithms that employs entropy regularization to stabilize policy gradient–based learning processes. However, it still exhibits sensitivity to parameter tuning, which can lead to unstable convergence performance or even divergence during training [leahy2022convergence, schulman2017proximal, mnih2016asynchronous]. To translate the complete ranking of candidate reasoning chains into effective training signals, we adopt a linear reward shaping scheme. We then define the advantage value $a_{j}$ of candidate $c_{j}$ as:
@@ 5.1 Experimental Setting
Datasets. We evaluate our approach on two benchmark medical QA datasets. PubMedQA [jin2019pubmedqa] contains biomedical research questions derived from PubMed abstracts, requiring yes/no/maybe answers based on scientific evidence. MedQA-USMLE [jin2021disease] consists of multiple-choice diagnostic questions from the United States Medical Licensing Examination, testing clinical reasoning and diagnostic capabilities. These datasets represent complementary aspects of medical reasoning: research-oriented knowledge synthesis and clinical diagnostic decision-making.
In addition, we further evaluate our model on the real-world Far Eastern Memorial Hospital (FEMH) clinical dataset, which consists of de-identified electronic medical records and physician diagnostic summaries. Unlike benchmark QA datasets, FEMH contains free-text clinical presentations and complex multi-symptom cases, providing a more realistic assessment of clinical reasoning performance.
Implementation Details. Our base model is Gemma 2B, chosen for computational efficiency while maintaining competitive performance. We generate five CoT reasoning variants per question and select the four highest-ranking candidates based on their acceptance probabilities for RPRO training. The acceptance threshold $\tau$ is set to 0.6 based on preliminary experiments. All models are trained with AdamW optimizer using a learning rate of 5e-5, batch size of 16, and 3 training epochs. Training is performed on NVIDIA A100 GPUs with mixed precision.
@@ 5.2 Baselines
We compare against a diverse set of established language models spanning different parameter scales, training strategies, and levels of medical specialization. General-purpose LLMs include Gemma 7B [mesnard2024gemma], Mistral 7B [jiang2023mistral], Falcon 7B [penedo2023falcon], Qwen2.5 7B, and Vicuna 7B. These models represent strong modern baselines trained on broad web-scale corpora.
For medical-specialized models, we evaluate MedAlpaca 7B and 13B [han2023medalpaca], which are fine-tuned specifically for medical Q&A tasks, and BioMistral 7B [labrak2024biomistral], trained on biomedical literature and clinical text. We additionally include Medicine-LLM, a medical-domain model tailored for clinical reasoning tasks, and Meditron 7B [chen2023meditron], a strong medical adaptation of LLaMA trained on PubMed and clinical corpora. Finally, GPT-OSS 20B[agarwal2025gpt] is included as a large open-source frontier model with strong general and biomedical reasoning performance.
Together, these baselines cover a wide spectrum of model sizes 7B–20B), training regimes (general-purpose vs. biomedical-pretrained vs. medically fine-tuned), and architectural families, offering a comprehensive comparison against both domain-general and domain-specific state-of-the-art systems.
@@ 5.4 Evaluation Metrics
Table 1: Comparison of different models on PubMedQA and MedQA-USMLE under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
Models PubMedQA (Accuracy / Macro F1) MedQA-USMLE (Accuracy / Macro F1)
0-shot 1-shot 5-shot 0-shot 1-shot 5-shot
BioMedLM 35.10 / 20.45 42.23 / 25.75 48.65 / 30.30 22.68 / 13.73 26.07 / 15.92 31.18 / 20.36
BioMistral 7B 36.89 / 27.98 42.76 / 32.41 45.53 / 36.28 19.89 / 7.47 31.53 / 13.41 37.57 / 16.10
Falcon 7B 55.83 / 37.72 57.14 / 39.08 57.92 / 40.63 17.23 / 8.14 24.87 / 16.21 24.89 / 13.94
Gemma 7B 58.08 / 48.02 58.67 / 49.51 58.12 / 50.84 20.37 / 10.18 43.12 / 24.15 51.18 / 32.35
MedAlpaca 7B 44.83 / 42.41 46.72 / 43.94 48.51 / 45.34 24.01 / 16.37 41.27 / 31.46 47.82 / 36.38
Meditron 7B 55.32 / 44.82 58.65 / 43.50 60.44 / 46.27 19.23 / 14.77 27.88 / 19.94 34.01 / 24.33
Medicine-LLM 41.87 / 33.76 42.25 / 35.91 42.18 / 35.58 26.35 / 16.90 34.37 / 23.46 39.82 / 28.47
Mistral 7B 51.43 / 47.11 54.03 / 49.81 56.78 / 52.43 22.66 / 11.78 40.85 / 23.47 49.02 / 29.16
Qwen2.5 7B 51.34 / 48.40 54.43 / 46.82 55.38 / 43.44 35.06 / 23.84 44.23 / 31.02 50.87 / 37.68
Vicuna 7B 56.27 / 49.13 58.88 / 50.34 59.92 / 53.47 27.43 / 17.86 33.91 / 22.56 40.12 / 27.34
MedAlpaca 13B 52.61 / 46.16 52.47 / 47.93 54.38 / 49.29 23.15 / 22.42 39.36 / 25.27 45.31 / 26.54
GPT-OSS 20B 47.27 / 43.75 51.23 / 47.95 55.61 / 52.34 40.02 / 23.76 50.13 / 30.42 52.68 / 32.11
Our Method (Gemma 2B, RPRO) 60.23 / 51.78 61.76 / 50.95 62.02 / 53.12 42.38 / 30.51 46.27 / 32.68 51.67 / 35.24
Table 2: Comparison of different models on the Far Eastern Memorial Hospital (FEMH) clinical dataset using BERTScore-F1 and Cosine Similarity under 0-shot, 1-shot, and 5-shot settings. Best results are in bold.
Models 0-shot (BERTScore-F1 / Cosine) 1-shot (BERTScore-F1 / Cosine) 5-shot (BERTScore-F1 / Cosine)
BioMedLM 0.798 / 0.342 0.812 / 0.351 0.825 / 0.353
BioMistral 7B 0.830 / 0.398 0.842 / 0.326 0.885 / 0.405
Falcon 7B 0.856 / 0.342 0.861 / 0.351 0.872 / 0.368
Gemma 7B 0.871 / 0.367 0.874 / 0.381 0.868 / 0.373
MedAlpaca 7B 0.848 / 0.334 0.852 / 0.343 0.861 / 0.357
Mistral 7B 0.859 / 0.345 0.865 / 0.359 0.871 / 0.366
Vicuna 7B 0.872 / 0.358 0.877 / 0.369 0.884 / 0.391
GPT-OSS 20B 0.851 / 0.338 0.864 / 0.354 0.873 / 0.368
Our Method (Gemma 2B, RPRO) 0.879 / 0.423 0.884 / 0.507 0.891 / 0.528
Table 3: Ablation study comparing different training methods and refinement strategies on PubMedQA and MedQA-USMLE under zero-shot settings. Best results are in bold..
@@
SFT (with Refinement) 57.43 40.53 37.08 23.72
DPO (with Refinement) 59.90 42.72 38.26 24.35
RPRO (No Refinement) 59.29 37.41 37.53 23.47
Our Method: RPRO (with Refinement) 60.23 51.78 42.38 30.51
Table 4: Reasoning results from FEMH clinical dataset
[…]
Input (Symptoms/Objective):
S: 1101201 Enzyme decrease $\rightarrow$ possible improved myocarditis… 110/11/26 2nd Moderna $\rightarrow$ 11/27 fever $\rightarrow$ 11/29 mild dyspnea, chest tightness (worse when lying) … LMD CPK WNL, TnT 1559 → 210. EKG: mild pericarditis (r/o). CXR: normal heart size.
O: BP, HR stable; clear BS; soft abdomen; no edema.
Model Reasoning: Step 1. Identify the key symptoms: The patient presents with fever, dyspnea, chest tightness, and mild discomfort. Step 2. Consider the laboratory results: The patient has elevated troponin levels, but CK-MB is normal. This suggests a possible myocarditis. Step 3. Integrate the clinical findings: The patient’s symptoms and laboratory results suggest a possible myocarditis. The EKG is normal, and the CXR is normal. Step 4. Consider differential diagnoses: The differential diagnoses include myocarditis, pericarditis, and other causes of chest pain and dyspnea. Diagnosis: Myocarditis. Diagnosis: Myocarditis. Prediction & Answer:
Prediction: Myocarditis.
True Diagnosis: r/o mild myocarditis.
[…]
Input (Symptoms/Objective):
S: Stinging sensation (2021/11), high IOP (>>50), blurred vision OD. Lost f/u 2 years. O: IOP: 53→17 (treated). Cornea clear. PCIOL subluxation inferiorly. OS: NS 1+–2+. Fundus: C/D 0.5–0.6 (OD). OCT RNFL OD 32.
Model Reasoning: Step 1. Potential causes: corneal opacity, cataract, glaucoma, IOL subluxation. Step 2. History: high IOP, decreased vision, recent IOL subluxation. Step 3. Comprehensive exam to assess cornea/lens/retina. Step 4. Conclusion: likely IOL subluxation related to prior surgery/high IOP. Diagnosis: IOL subluxation. Diagnosis: IOL subluxation. Prediction & Answer:
Prediction: IOL subluxation.
True Diagnosis: IOL dislocation (OD).
[…]
Input (Symptoms/Objective):
S: Pain over lower right gingiva during upper RPD delivery.
O: Full-mouth chronic periodontitis, Upper/Lower RPD mostly acceptable except #46 supraocclusion, Ulceration on lower right edentulous ridge
Model Reasoning: Step 1. Assess the presence of an ulcer on the lower right edentulous ridge. Step 2. Determine if there is any retention or stability issue with the RPDs. Step 3. Evaluate the condition of the gingiva in relation to the RPDs. Step 4. Consider the history of chronic periodontal disease. Diagnosis: Chronic periodontal disease with ulceration. Prediction & Answer:
Prediction: Chronic periodontal disease with ulceration.
True Diagnosis: (1) Full-mouth chronic periodontitis. (2) Ulceration over lower right edentulous ridge mucosa.